First Draft of What I Need for conversion e^{+}e^{−} analysis M. J. Tannenbaum Electron WG 10/97

For Each Event

• List of electron candidates with Very Loose Cuts

• Event Vertex Global Information to Characterize Reaction

• E_{T} (η) in e.g. 7 bins of δη = 0.1 (−0.35 ≤ η ≤ −0.35) from EMCal • dn(η)/dη from MVD

a. in central acceptance b. in full MVD acceptance

• Zero Degree Cal Energy from ZDC

• Number, <p_{T} > or <m_{T} − m> for Charged Tracks a. All detected non-identified charged particles b. All identified π^{+}, π^{−}, K^{+}, K^{−}, p

• <E_{T} > and number of clusters for several classes EMCal a. All calorimter clusters b. All calorimeter clusters caused by charged track c. All non-charged track calorimeter clusters d. All calorimeter clusters E_{T} ≥ 500 MeV e. ...

M. J. Tannenbaum, PHENIX Electron WG–Requirements for Conv. 10/97 2

Analysis Algorithm

1. Histogram p_{T} of electrons with loose cuts, + and − separately (will use to find Efficiency of cuts)

2. Histogram p_{T} of electrons with best cuts, + and − separately

3. Start with first member of list

4. Loop over all other members of list looking for a. pairs close in θ : Δθ ≤ ‘Pair cut’ (Conversion + Dalitz) b. pairs somewhat close in θ: ‘Pair cut’ ≤ Δθ ≤ ‘Dalitz cut’

(Dalitz + ‘Direct Photon’ + ) c. Alternatively Loop over all pairs with no θ cut— (Combanitorics!) d. Figure out how to reduce combinatorics, e.g. by using n small regions

of size ΔΩ/n, for instance slices in η for high rate high background processes like low mass pairs which are very collimated in ϕ, η before magnet but open up in ϕ after magnet.

5. For Each Pair calculate:

a. 4-vector b. pair vertex c. m, p_{T}, δm, δp_{T} (resolutions)

6. Histogram pair vertex [r, ϕ, Z]

7. Separate, if possible, pairs with vertex a. In beam region b. Near known converters e.g. beam pipe, MVD ... c. elsewhere

8. Plot for ‘clear’ conversions in Beam Pipe and MVD:

\frac{dσ}{p_{T} dp_{T} dη} also \frac{dσ}{dm}

in p_{T}, m bins—probably 2-d weighted histogram. needs some thought

M. J. Tannenbaum, PHENIX Electron WG–Requirements for Conv. 10/97 3

9. Ditto for ‘clear’ dalitz from Beam Line 8a,9a Separate plots for ‘Peripheral’, ‘non-Peripheral’ and ‘Central’ events 8b,9b Plot same charge vs opposite charge for Actual Events 8c,9c Plot same charge vs opposite charge for Mixed Events

10. Derive inclusive γ spectrum and π^{0} spectrum from 8., 9. and check whether they are consistent

* Analysis Needed to figure out how to do this

11. Check conversion measurement by making dedicated runs:

a. With extra %’s radiation lengths converter added near MVD

Now that all the hard work is done—Cash In

12. Plot e^{±}/{ Dalitz (+ Conversion) e^{+}e^{−}} vs p_{T}

* There is a simple-to-calculate curve, probably a constant, if all e^{±} due to Dalitz + conversions

The data will deviate from this curve for p_{T} ≥ 1 GeV/c, allowing us to clearly see a ‘charm’ signal, probably e^{±} for p_{T} ≥ 2.0 Gev/c.

13. Plot invariant masses of All same charge and opposite charge pairs 14. Due to PHENIX excellent mass resolution, find J/Ψ peak if luminosity

is adequate, and if histogram binning is reasonable.

15. Plot Ψ/{e^{±} > 2.0 GeV/c} ∝ Ψ/c − c̄ vs E_{T}

16. Find ϕ peak and plot ϕ/dalitz vs. p_{T}, ϕ/c − c̄ vs E_{T} ...

17. Look at other mass regions and their ratios a. Ψ/Drell-Yan b. ϕ/Drell-Yan c. ... Plot these ratios vs E_{T} à la Gonin

M. J. Tannenbaum, PHENIX Electron WG–Requirements for Conv. 10/97 4

18. Go to a. Stockholm (Nobel Prize) b. Lund (Collaboration Meeting) Depending on the result!!!